OpenStudy (ksaimouli):

4-5e^3x=-26 solve the exponenitial function

OpenStudy (ksaimouli):

help

OpenStudy (mertsj):

Subtract 4 from both sides and then divide both sides by -5

OpenStudy (mertsj):

Then take the natural log of both sides.

OpenStudy (rogue):

\[4-5e^{3x}=-26\]\[30 = 5 e^{3x}\]\[e^{3x} = 6\]\[\ln \left[ e^{3x} \right] = \ln 6\]\[3x = \ln 6\]\[x = \frac {\ln 6}{3}\]

OpenStudy (ksaimouli):

\[4-5e^3x=-26\]